The Loftis Company is preparing its pro forma financial statements for the next year using this model. The abbreviated financial statements are presented below: Sales Growth 20% Tax Rate 34% Income Statement Sales $780,000 Costs 415,000 Depreciation 135,000 Interest 68,000 Taxable Income $162,000 Taxes 55,080 Net Income $106,920 Dividends 30,000 Additional Retained Earnings $76,920 Balance Sheet Assets Liabilities and Owner's Equity Current Assets $240,000 Total Debt $880,000 Net Fixed $1,350,000 Owners Equity $710,000 $1,590,000 $1,590,000 Required: A. Calculate the Parameter Estimates Cost Percentage = (Cost/Sales) Depreciation Rate = (Depreciation/Beginning Fixed Assets) Interest Rates = (interest paid/total debt) Tax Rate = (taxes/net income) Payout ratio = (dividends/ net income) Capital Intensity Ratio = (Fixed Assets /Sales) REQUIRED: b. Construct the pro forma financial statements using the parameters you calculated. Your pro forma balance sheet should balance. Loftis Company "Pro forma Income Statement" Sales $- Costs - Depreciation - Interest - Taxable income $- Taxes - Net income $- Dividends Additions to retained earnings Loftis Company "Pro forma Balance Sheet" Assets Liabilities and Equity Current assets $- Total debt $- Net fixed assets - Owners' equity - Total assets $- Total debt and equity $- b. Construct the pro forma balance sheet. What is the total debt necessary to balance the pro forma balance sheet? c. Explain why and/or when you would "implement" this approach in an organization (Short Essay 125 to 150 words) d. Discuss several means of determining the effectiveness of the "outcomes" or results using this approach (Short Essay 125 to 150 words)

I would like to not only see the answer but the logic of finding the answer
Answer #1

Part (a)

Please see the table below. In the last part, all the percentages gave been calculated. Second column titled "Linkage" explains how each row has been calculated.

Parameter Linkage $
Sales A        780,000
[-] Costs B        415,000
[-] Depreciation C        135,000
[-] Interest D          68,000
Taxable Income E        162,000
[-] Taxes F          55,080
Net income G        106,920
[-] Dividends H          30,000
Addition to retained earnings I          76,920
Current Assets P        240,000
[+] Net Fixed Assets Q     1,350,000
Total Assets R     1,590,000
Total Debt S        880,000
[+] Owner's Equity T        710,000
Total debt & equity U     1,590,000
Beginning fixed assets V = Q + C     1,485,000
A. Cost percentage CP = B / A 53.21%
B. Depreciation Rate DR = C / V 9.09%
C. Interest rate IR = D / S 7.73%
D. Tax Rate TR = F / E 34.00%
E. Payout ratio PR = H / G 28.06%
F. Capital Intensity Ratio CR = Q / A 1.73

Part (b)

Please see the tables below for proforma. Please see the last column for the calculation behind the value.

Parameter $ How it has been calculated?
Sales        780,000 780000 x (1 + 20%)
[-] Costs        415,000 53.21% x 936000
[-] Depreciation        135,000 9.09% x 1350000
[-] Interest          68,000 7.73% x 880000
Taxable Income        162,000 Sales - all the cost above
[-] Taxes          55,080 34.00% x 247273
Net income        106,920 247273 - 84073
[-] Dividends          30,000 28.06% x 163200
Addition to retained earnings          76,920 163200 - 45791
Current Assets        240,000 240000 x (1 + 20%)
[+] Net Fixed Assets     1,350,000 936000 x 1.73
Total Assets     1,590,000 sum of the two items above
Total Debt        880,000 1908000-827409
[+] Owner's Equity        710,000 710000+76920
Total debt & equity     1,590,000 sum of the two items above

The total debt necessary to balance the pro forma balance sheet = Total assets - owner's equity = 1,908,000-827,409 = 880,000

Part (c)

(Please build upon these points to get the desired length of 100 - 150 words)

This approach would be implemented by the organization:

  • When the business model is simple and can be expresses as %age of sales
  • Cost structure is mostly variable and is linked to sales fully. If the company has a higher proportion of fixed costs in the operations, this method will project higher costs than actual.
  • Debt is constant and no repayment is made during the year
  • All the assets are spontaneous that is they rise in proportion to sales
  • No fresh infusion of equity is envisaged.

Part (d)

(Please build upon these points to get the desired length of 100 - 150 words)

Means to examine the effectiveness of this method:

  • As a first cut, this method is effective in figuring out the dependence on external capital for the next year or period
  • It tells you how much is the internal accruals going to be.
  • We can see the nature of our costs and segregate them into fixed and variable. That will help us to improve the effectiveness of this forecast.
  • Results of this model can tell a firm in advance whether there will be operational shortfall or surplus and hence act as early warning for the firms to start preparing for fund raise.